OpenStudy (kirbykirby):

How are these equal? \[\lim_{x \rightarrow \infty}\frac{\sin(1/x)}{1/x}=\lim_{\theta \rightarrow 0}\frac{\sin \theta}{\theta}=1\]

OpenStudy (kirbykirby):

I know that lim sinx/x as x-> 0 is 1, but im not sure how the leftmost part becomes equal to the right part

OpenStudy (zarkon):

let \[\theta=\frac{1}{x}\]

OpenStudy (kirbykirby):

how does the limit go from infinity to 0?

OpenStudy (kirbykirby):

tha's the part i'm having trouble with ._.

OpenStudy (zarkon):

if \[x\to \infty\text{ then }\frac{1}{x}\to 0\]

OpenStudy (kirbykirby):

ohh.. right =_= omg I feel so dumb lol. Thanks :D